# Ludax: A GPU-Accelerated Domain Specific Language for Board Games

### v1

@@ 5 Expressive Range
As mentioned above, Ludax currently supports a relatively narrow class of games: two-player, perfect-information board games played by placing, capturing, and flipping pieces. Both Ludii and PGX contain many games that Ludax does not: PGX includes implementations of Backgammon, Chess, Shogi, and Go, along with hidden-information and Atari-style games (though Ludax also supports games that do not appear in PGX), while Ludii’s description language can encode a vast array of board game styles and subgenres. Despite this, Ludax’s description language remains quite expressive. In addition to simple $m-n-k$ line completion games, Ludax supports complex and asymmetric winning conditions (e.g. misère variants, score-based victory), piece capturing and flipping, directional adjacency checks and restrictions, “custodial” mechanics, and games based on connecting arbitrary board regions. Ludax also supports regular rectangular and hexagonal boards of arbitrary sizes, as well as “hexagonal-rectangular” boards (e.g. as used in Hex). These components can then be combined compositionally to form a wide array of unique mechanics and dynamics. In addition, because Ludax is a general description language, implementing a single new game component expands the entire space of games in the framework. While the class of games representable in Ludax may at present be smaller than that of Ludii or other game description languages, it remains expansive.
@@
6 Benchmarking
We benchmark the speed of Ludax on a set of 8 games, 4 of which are also implemented in both Ludii and PGX (allowing for a full comparison) and 4 of which are implemented only in Ludax and Ludii. Again, we emphasize that these 8 games are just exemplars of the class of games which Ludax supports, not an exhaustive list. A full description of each benchmark game is available in the Supplementary Material. We perform each of our benchmarking experiments on a workstation with a single NVIDIA 4090 GPU, 32 CPU cores, and 128GB of memory. In Figure 2 we plot the throughput (in steps per second) under a uniformly random action policy for each game environment against the batch size (log scale on both axes), with the standard deviation of throughputs across episodes as error bars. Ludii supports parallelization via multi-threading: we report throughput on the same workstation when parallelized on 1, 16, and 32 threads. Evaluations for Ludax and PGX were obtained by performing 100 warmup full-game episodes at the specified batch size, followed by measuring the speed over 500 episodes, with each evaluation taking at most a few minutes to complete. Evaluations for Ludii were obtained by running warmup episodes for 10 seconds, followed by measuring the speed over 30 seconds of episodes. [^2] We opted to measure speed for Ludax and PGX using a fixed number of episodes because JAX’s compilation procedure makes it difficult to halt execution after a specific elapsed wall time.
Overall, Ludax achieves speeds that are competitive with state-of-the-art JAX environments. At small batch sizes, its throughput is similar to that of the PGX implementations. At larger batch sizes in more complicated games (i.e. Hex and Reversi), PGX takes a clear edge – though Ludax remains within an order of magnitude of PGX. The comparative “plateauing” of Ludax’s speed at high batch sizes may be due to memory pressure – for instance, Ludax’s implementation of Hex maintains both a board and the connected components for each game state, whereas the PGX implementation cleverly combines both into a single array. This kind of optimization is of course theoretically implementable in Ludax as well, though again we emphasize the desiderata of avoiding game-specific solutions.
Ludax also outspeeds Ludii on 16 and 32 threads across all 8 games, achieving a maximum speedup of between $\sim$3x (Hex) and $\sim$55x (Pente). We note that there are factors that both advantage and disadvantage Ludax in this specific comparison against Ludii. One potential advantage for Ludax is its smaller representation space – implementations of basic mechanics in Ludii support a wider range of optional arguments and board types, with a corresponding increase in computational overhead (though see Section 4 for how this may be avoided). Conversely, Ludii’s ability to use dynamically-sized data structures brings advantages that are particularly beneficial in uniformly random playouts, but would (partially) disappear in playouts using deep reinforcement learning. Specifically, where Ludax samples random actions from categorical probability distribution by masking over the entire board at each step, Ludii selects random actions simply by taking a random index from a list. Ludii also has optimized playout implementations tailored towards the many of the categories of games covered by Ludax Soemers_2022_Optimised , though these optimizations are also more difficult to apply in the context of deep learning.
7 Learned Agents
@@
Finally, we demonstrate the feasibility of training reinforcement learning agents using the Ludax framework. We train our agent on the game Reversi (also known as Othello) using the AlphaZero-style silver2017mastering training script from the PGX library [^3] https://github.com/sotetsuk/pgx/blob/main/examples/alphazero/train.py (used under Apache 2.0 license) (making only slight modifications to accommodate minor differences between the Ludax and PGX APIs). We use the same ResNetV2 kaiming2016resnet network architecture and training hyperparameters as PGX (full details available in the Supplementary Material) and train three separate runs on a single A100 GPU. Each run lasted roughly 57 million frames and took roughly three hours to complete.
@@
We compare the performance of agents trained in the Ludax and PGX environments against the baseline Reversi agent provided by the PGX library in Figure 3. Evaluations were performed by playing two batches of 1024 games (one with the learned agent as the first player and one as the second player), with actions sampled from the normalized output of the policy head at each step. We see that both learned agents achieve remarkably similar performances against the baseline, with little to no differences in learning speed or stability. While a more thorough, tournament-based evaluation would be necessary to properly rank the agents against each other, our objective is to demonstrate the general success of the training procedure and not to definitively defeat the baseline agent. Although the PGX implementation of the Reversi environment is slightly more efficient, this translated into only marginal improvements in overall runtime owing to the shared overhead of network forward passes and weight updates. Like PGX, Ludax offers a familiar API and an efficient set of implementations with which to train learned player agents.
8 Limitations
@@
Generality: As mentioned in Section 5, Ludax currently supports a smaller class of games than other comparable game description languages. While we aim to increase the range of games expressible in Ludax (see below), it will likely never match the full generality of Ludii. As such, other frameworks may be more appropriate for use-cases in which a broad range of games is more important than rapid simulation. Further, Ludax does not support genres other than board games (e.g. video games, card games, …) – we leave the development of hardware accelerated description languages for such domains as an exciting area of future work.
Efficiency: Compared to bespoke JAX implementations of board games (such as in the PGX library), environments in Ludax have worse throughput – especially at larger batch sizes. While we deploy a number of optimizations to help close the efficiency gap when possible (see Section 4), there are ultimately unavoidable trade-offs between speed and generality. For the purpose of training or benchmarking single-task agents on existing games, hard-coded simulators are likely the superior choice. We note, however, that Ludax also provides a way for non-experts to design and benchmark on novel games without having to write any JAX code themselves.
9 Future Work
The most obvious avenue of extension for Ludax is the implementation of additional game mechanics. In particular, we aim to support games based on piece movement (e.g. Amazons), games with multiple piece types (e.g. Checkers), and games with multiple distinct gameplay phases (e.g. Nine-Men’s Morris). Other generalizations, such as support for irregular board shapes, would also help expand the space of games representable in Ludax. In addition, it’s also very likely that the implementation of specific gameplay elements could be further optimized for throughput and / or memory footprint. However, a balance must be struck between efficiency and generality: a less efficient solution which accommodates all valid games under the grammar is ultimately preferable to one which only applies to a subset of games. Lastly, we aim to provide a more robust visual interface for Ludax, both for the purpose of facilitating human-subject research and the potential development of more “human-like” artificial agents which process the game board at the pixel level and select actions spatially.
@@ 9 Future Work
We are particularly excited about the potential application of Ludax to the study of automated game design (or reward-guided program synthesis more generally Cui_2021_AlphaEvolve surina2025algorithmdiscoveryllmsevolutionary RomeraParedes2024FunSearch ). Systems like GAVEL Todd_2024_GAVEL depend on both a broad representation space and rapid evaluation of novel games. The efficiency of Ludax may make it possible to train a reinforcement learning agent from scratch as part of the inner loop of game evaluation, potentially unlocking a new range of computational features (e.g. learning curves) that correlate with human notions of fun and engagement. Relatedly, Ludax may prove useful to research on human behavior and play. Recent work has explored heuristic-based computational models of human play on simple line completion games zhang2024people , and Ludax offers the possibility to both accelerate computation and broaden the domain to a wider class of games. Finally, Ludax offers an avenue to extend recent research in general game playing (e.g. with large language models schultz2024mastering ) by providing a wide base of efficient game implementations that can in turn be leveraged for tree search algorithms or training world models.
@@ 10 Conclusions
We introduce a novel framework for games research that combines the generality of game description languages with the efficiency of modern hardware-accelerated learning environments. Our framework, Ludax, represents a broad class of two-player board games and compiles directly into code in the JAX Python library. Games in Ludax achieve speeds that are competitive with hand-crafted JAX implementations and faster than the widely-used Ludii game description language, and Ludax environments can easily be deployed in existing pipelines for deep reinforcement learning. Our framework helps generalize and accelerate games research, with the potential to unlock entirely new kinds of agents and systems.
Broader Impact
This paper presents a general framework with the goal of advancing reinforcement learning and games research. While there are many potential societal consequences of such work in general, we do not feel that any must be specifically highlighted here.
Appendix A Ludax Grammar
Below we present the complete grammar specification for Ludax, using the syntax of the Lark Python library.
game: "(game" name players equipment rules ")"
@@ Appendix A Ludax Grammar
players: "(players" positive_int ")"
@@
equipment: "(equipment" board ")"
board: "(board" (board_square | board_rectangle | board_hexagon | board_hex_rectangle) ")"
board_square: "(square" number ")"
board_rectangle: "(rectangle" number number ")"
board_hexagon: "(hexagon" number ")"
board_hex_rectangle: "(hex_rectangle" number number ")"
@@
rules: "(rules" start_rules? play_rules end_rules ")"
@@
start_rules: "(start" start_rule+ ")"
@@
start_place: "(place" player_reference pattern_arg ")"
@@
play_rules: "(play" play_phase+ ")"
@@
phase_once_through: "(once-through" play_mover_order play_mechanic ")"
phase_repeat: "(repeat" play_mover_order play_mechanic ")"
@@
play_mechanic: play_place force_pass?
play_place: "(place" mover_reference? place_destination_constraint place_result_constraint? play_effects?")"
force_pass: "(force_pass" ")"
// ---Constraints---
place_destination_constraint: "(destination" super_mask ")"
place_result_constraint: "(result" super_predicate ")"
@@
play_effects: "(effects" play_effect+ ")"
@@
effect_capture: "(capture" super_mask mover_reference? increment_score_arg? ")"
effect_flip: "(flip" super_mask mover_reference? ")"
effect_increment_score: "(increment_score" mover_reference function ")"
effect_set_score: "(set_score" mover_reference function ")"
@@
function_add: "(add" function+ ")"
function_connected: "(connected" multi_mask_arg mover_reference? direction_arg? ")"
@@
function_count: "(count" super_mask ")"
function_line: "(line" positive_int orientation_arg? exact_arg? ")"
function_multiply: "(multiply" function+ ")"
function_score: "(score" mover_reference ")"
function_subtract: "(subtract" function function ")"
@@
end_rules: "(end" end_rule+ ")"
end_rule: "(if" super_predicate end_rule_result ")"
@@
result_win: "(" mover_reference "win" ")"
result_lose: "(" mover_reference "lose" ")"
@@
super_mask: mask | super_mask_and | super_mask_or | super_mask_not
super_mask_and: "(and" super_mask+ ")"
super_mask_or: "(or" super_mask+ ")"
super_mask_not: "(not" super_mask ")"
@@
mask_adjacent: "(adjacent" super_mask direction_arg? ")"
mask_center: "center"
mask_corners: "corners"
mask_custodial: "(custodial" custodial_length_arg mover_reference? orientation_arg? ")"
mask_edge: "(edge" edge ")"
mask_empty: "empty"
mask_occupied: "occupied" | "(occupied" mover_reference ")"
mask_pattern: "(pattern" dimensions_arg pattern_arg rotate_arg? ")"
mask_prev_move: "(prev_move" mover_reference ")"
@@
multi_mask_corners: "corners"
multi_mask_edges: "edges"
multi_mask_edges_no_corners: "edgesNoCorners"
@@
super_predicate_and: "(and" super_predicate+ ")"
super_predicate_or: "(or" super_predicate+ ")"
super_predicate_not: "(not" super_predicate ")"
predicate: predicate_equals
@@
predicate_equals: "(=" function+ ")"
predicate_exists: "(exists" super_mask ")" // technically equivalent to (>= (count mask) 1)
@@
predicate_greater_equals: "(>=" function function ")"
predicate_less_equals: "(<=" function function ")"
predicate_mover_is: "(mover_is" player_reference ")"
predicate_passed: "(passed" (mover_reference | BOTH) ")"
@@
dimensions_arg: "(" positive_int positive_int ")"
direction_arg: "direction:" direction
@@
multi_mask_arg: multi_mask | "(" super_mask+ ")"
@@
pattern_arg: "(" positive_int+ ")"
@@
?direction: UP | DOWN | LEFT | RIGHT | UP_LEFT | UP_RIGHT | DOWN_LEFT | DOWN_RIGHT | VERTICAL | HORIZONTAL | ORTHOGONAL | DIAGONAL | BACK_DIAGONAL | FORWARD_DIAGONAL | ANY
@@
// Constants
TOP: "top"
BOTTOM: "bottom"
UP: "up"
DOWN: "down"
LEFT: "left"
RIGHT: "right"
TOP_LEFT: "top_left"
TOP_RIGHT: "top_right"
BOTTOM_LEFT: "bottom_left"
BOTTOM_RIGHT: "bottom_right"
UP_LEFT: "up_left"
UP_RIGHT: "up_right"
DOWN_LEFT: "down_left"
DOWN_RIGHT: "down_right"
VERTICAL: "vertical"
HORIZONTAL: "horizontal"
ORTHOGONAL: "orthogonal"
DIAGONAL: "diagonal"
BACK_DIAGONAL: "back_diagonal"
FORWARD_DIAGONAL: "forward_diagonal"
ANY: "any"
TRUE: "true"
FALSE: "false"
MOVER: "mover"
OPPONENT: "opponent"
P1: "P1"
P2: "P2"
BOTH: "both"
Appendix B Benchmark Game Descriptions
Below, we present natural language descriptions of the rules for each of the exemplar games analyzed in Section 6.

### v2

@@ 5 Expressive Range
Ludax supports a class of two-player, perfect-information board games played by placing, capturing, and moving game pieces of different types (and can be adapated to support simple gridworlds – see Appendix D). Supported games range from the very simple (e.g. Tic-Tac-Toe) to the relatively complex (e.g. English Draughts, which features multiple kinds of pieces, forced captures, and sophisticated logic for granting extra turns). Existing games can also be easily modified, for instance by creating misère variants or modifying the size and shape of the board. More importantly, the low-level mechanics used to implement individual games can be combined compositionally to form a vast array of unique interactions and dynamics. Further, because Ludax is a general description language and not a catalog of games, adding a single new mechanic or component expands the entire space of representable games.
6 Benchmarking
We benchmark the speed of Ludax on a set of 10 games, 4 of which are also implemented in both Ludii and PGX (allowing for a full comparison) and 6 of which are implemented only in Ludax and Ludii. Again, we emphasize that these 10 games are just exemplars of the class of games that Ludax supports, not an exhaustive list. A full description of each benchmark game is available in the Supplementary Material. We perform each of our benchmarking experiments on a workstation with a single NVIDIA 4090 GPU and an Intel i9-9980XE processor (36 threads on 18 CPU cores and 128GB of memory).
In Figure 2 we plot the throughput (in steps per second) under a uniformly random action policy for each game environment against the batch size (log scale on both axes), with the standard deviation of throughputs across episodes as error bars. Ludii supports parallelization via multi-threading: we report the fastest throughput on the same workstation when parallelized on 1, 16, and 32 threads. We note, however, that Ludii does not natively support reinforcement learning with fixed-size observation and action spaces (see Soemers_2022_DeepLearning for a detailed discussion). To give a more realistic comparison in the context of RL specifically, we additionally report the throughput for Ludii when it constructs a fixed-size tensor for each state and indexes into a fixed-size approximation of the action space for each legal action. Evaluations for Ludax and PGX were obtained by performing 100 warmup full-game episodes at the specified batch size, followed by measuring the speed over 500 episodes, with each evaluation taking at most a few minutes to complete. Evaluations for Ludii were obtained by running warmup episodes for 10 seconds, followed by measuring the speed over 30 seconds of episodes. [^1] We opted to measure speed for Ludax and PGX using a fixed number of episodes because JAX’s compilation procedure makes it difficult to halt execution after a specific elapsed wall time. For games with potentially unbounded length (e.g. Dai Hasami Shogi), we terminate games for both Ludax and Ludii after 200 total turns.
Overall, Ludax achieves speeds that are competitive with state-of-the-art JAX environments optimized for a single game. PGX takes a slight edge in throughput at large batch sizes for Connect Four and Reversi, but Ludax remains within an order of magnitude (and even outspeeds PGX in Hex [^2] Though Ludax’s and PGX’s implementations differ slightly with the latter including the “swap rule,” this should not have a large effect on throughput under random playouts). Ludax also outspeeds Ludii across all 10 games, achieving a maximum speedup of between 3x (English Draughts) and 450x (Hex) in the tensor playout regime. Some of this speedup is likely due to Ludax’s comparatively smaller representation space, though Ludii also has optimized playout implementations tailored towards many of the categories of games covered by Ludax which might be difficult to apply in the context of deep learning (Soemers_2022_Optimised).
7 Trained Agents
@@
Finally, we demonstrate the feasibility of training reinforcement learning agents using the Ludax framework. We train our agent on the game Reversi (also known as Othello) using the AlphaZero-style (silver2017mastering) training script from the PGX library [^3] https://github.com/sotetsuk/pgx/blob/main/examples/alphazero/train.py (used under Apache 2.0 license). We use the same ResNetV2 (kaiming2016resnet) network architecture and training hyperparameters as PGX (full details available in the Supplementary Material) and train three separate runs on a single A100 GPU. Each run lasted roughly 57 million frames and took roughly three hours to complete.
@@
We compare the performance of agents trained in the Ludax and PGX environments against the baseline Reversi agent provided by the PGX library in Figure 3. Evaluations were performed by playing two batches of 1024 games (one with the learned agent as the first player and one as the second player), with actions sampled from the normalized output of the policy head at each step. We see that both learned agents achieve remarkably similar performances against the baseline, with little to no differences in learning speed or stability. Although the PGX implementation of the Reversi environment is slightly more efficient, this translated into only marginal improvements in overall runtime (about $1.5\%$) owing to the shared overhead of network forward passes and weight updates. We also report training runs for Hex and Connect Four in Appendix F. Like PGX, Ludax offers a familiar API and an efficient set of implementations with which to train learned player agents.
8 Limitations and Future Work
@@
While we aim to continually expand the range of games expressible in Ludax, it will likely never match the full generality of a language like Ludii. As such, other frameworks may be more appropriate for use-cases in which a broad range of games is more important than rapid simulation. Ludax also does not support genres other than board games (e.g. video games, card games, …) – we leave the development of hardware accelerated description languages for such domains as an exciting area of future work. Compared to bespoke JAX implementations of board games (such as in the PGX library), environments in Ludax have slightly worse throughput – though the gap is marginal in a standard RL training setup. We deploy a number of optimizations to help close the efficiency gap when possible (see Section˜4), but there are ultimately unavoidable trade-offs between speed and generality. For the purpose of training or benchmarking single-task agents on existing games, hard-coded simulators may remain the superior choice.
In light of these limitations, the most obvious avenues of extension for Ludax are the implementation and optimization of additional games and mechanics mechanics like irregular board shapes (e.g. Tabu Y) and mechanically distinct gameplay phases (e.g. Nine-Men’s Morris). We also aim to provide a more robust visual interface for Ludax, both for the purpose of facilitating human-subject research (carvalho2025nicewebrl) and the potential development of more “human-like” artificial agents which process the game board visually and select actions spatially.
@@
Looking further ahead, we are excited about the potential application of Ludax to the study of automatic game design. Such systems depend on both a broad representation space and rapid evaluation of novel games – see Appendix G for a preliminary investigation of Ludax’s suitability for this kind of research. Ludax may also prove useful to research on human behavior and play, for instance by expanding and accelerating work that models human player behavior (zhang2024people). Finally, Ludax offers an avenue to extend recent research in general game playing (e.g. with large language models (schultz2024mastering)) by providing a wide base of efficient game implementations that can in turn be leveraged for tree search algorithms or training world models.
@@ 9 Conclusions
We introduce a novel framework for games research that combines the generality of game description languages with the efficiency of modern hardware-accelerated learning environments. Our framework, Ludax, represents a broad class of two-player board games and compiles directly into code in the JAX Python library. Games in Ludax achieve speeds that are competitive with hand-crafted JAX implementations and faster than the widely-used Ludii game description language, and Ludax environments can easily be deployed in existing pipelines for deep reinforcement learning. Our framework helps widen and accelerate games research, with the potential to unlock new approaches in RL generalization, automatic game generation, and cognitive modeling.
Appendix A Example Games and Syntax
In Figure 4 we present the Ludax syntax for a small set of exemplar games (Reversi, Connect Four, Yavalath, and Wolf and Sheep). Below, we present an annotated version of English Draughts to help illustrate aspects of Ludax’s syntax and structure.
(game "English Draughts"
// We define the "forward" direction for each player to simplify the movement logic
(players 2
(set_forward (P1up) (P2down))
// The game takes place on an 8-by-8 board and uses two kinds of pieces
(equipment
(board (square 8))
(pieces
("pawn" both)
("king" both)
(rules
// To begin, we place pawns for each player in a checkerboard pattern along the first and last two rows
(start
(place "pawn" P1 (40 42 44 46 49 51 53 55 56 58 60 62))
(place "pawn" P2 (1 3 5 7 8 10 12 14 17 19 21 23))
(play
// Players alternate making moves
(repeat (P1 P2)
// A move consists of one of the following options:
(move
(or
// A pawn can hop over opposing pieces in either of the "forward" diagonals to capture them. This option has priority over non-capture moves
(hop "pawn" direction:(forward_left forward_right) hop_over:opponent capture:true priority:0)
// A pawn can step into unoccupied squares in the "forward" diagonals
(step "pawn" direction:(forward_left forward_right) priority:1)
// A king hops and captures in any diagonal direction...
(hop "king" direction:diagonal hop_over:opponent capture:true priority:0)
// ...and steps the same way
(step "king" direction:diagonal priority:1)
(effects
// After moving, pawns on the forward edge of the board are promoted into kings
(promote "pawn" "king" (edge forward))
// If the player made a hop move (i.e. captured a piece) and the same piece could hop and capture again, the player gets an extra turn using that piece
(if (and (action_was mover hop) (can_move_again hop))
(extra_turn mover same_piece:true)
// If making a move results in the next player having no legal moves, then the most recent mover wins
(end
(if (no_legal_actions) (mover win))
// For rendering, player one has the black pieces and player two has the white pieces
(rendering
(color P1 black)
(color P2 white)
Figure 4: Ludax syntax for Reversi, Connect Four (classic board games), Yavalath and Wolf and Sheep. Ludax supports a wide range of games of vary complexity.
Appendix B Ludax Grammar
Below we present the complete grammar specification for Ludax, using the syntax of the Lark Python library (raw string constants omitted for brevity).
// ---Root---
game: "(" "game" name players equipment rules rendering? ")"
@@ Appendix B Ludax Grammar
players: "(" "players" positive_int forward_assignments? ")"
forward_assignments: "(" "set_forward" p1_assignment p2_assignment ")"
p1_assignment: "(" P1 (UP | DOWN | LEFT | RIGHT) ")"
p2_assignment: "(" P2 (UP | DOWN | LEFT | RIGHT) ")"
@@
equipment: "(" "equipment" board pieces regions? ")"
board: "(" "board" (board_square | board_rectangle | board_hexagon | board_hex_rectangle) ")"
?board_square: square_shape
?board_rectangle: rectangle_shape
?board_hexagon: hexagon_shape
?board_hex_rectangle: hex_rectangle_shape
// ---Pieces---
pieces: "(" "pieces" piece_definition+ ")"
piece_definition: "(" name (player_reference | BOTH) ")"
piece_reference: name
// ---Regions---
regions: "(" "regions" region_definition+ ")"
region_definition: "(" name (indices_arg | multi_mask_arg) ")"
region_reference: name
@@
rules: "(" "rules" start_rules? play_rules end_rules ")"
@@
start_rules: "(" "start" start_rule+ ")"
@@
start_place: "(" "place" piece_reference player_reference (indices_arg | multi_mask_arg) ")"
@@
play_rules: "(" "play" play_phase+ ")"
@@
phase_once_through: "(" "once_through" play_mover_order play_super_mechanic ")"
phase_repeat: "(" "repeat" play_mover_order play_super_mechanic ")"
@@
play_super_mechanic: play_mechanic force_pass?
play_mechanic: play_place | play_move
force_pass: "(" "force_pass" ")"
// ---Place rules---
play_place: "(" "place" piece_reference mover_reference? place_destination_constraint place_result_constraint? play_effects? ")"
place_destination_constraint: "(" "destination" super_mask ")"
place_result_constraint: "(" "result" super_predicate ")"
// ---Move rules---
play_move: "(" "move" move_type_definition play_effects? ")"
move_type_definition: move_type | "(" "or" move_type+ ")"
move_type: move_hop
| move_slide
| move_step
move_hop: "(" "hop" piece_reference direction_arg? piece_arg? hop_over_arg? capture_arg? priority_arg? ")"
move_slide: "(" "slide" piece_reference direction_arg? distance_arg? priority_arg? ")"
move_step: "(" "step" piece_reference direction_arg? priority_arg? ")"
@@
play_effects: "(" "effects" play_super_effect+ ")"
play_super_effect: play_effect | play_conditional_effect
play_conditional_effect: play_if_effect | play_if_else_effect
play_if_effect: "(" "if" super_predicate play_effect ")"
play_if_else_effect: "(" "if" super_predicate play_effect "else" play_effect ")"
@@
effect_capture: "(" "capture" super_mask mover_arg? increment_score_arg? ")"
effect_extra_turn: "(" "extra_turn" mover_reference same_piece_arg? ")"
effect_flip: "(" "flip" super_mask mover_arg? ")"
effect_increment_score: "(" "increment_score" mover_reference function ")"
effect_promote: "(" "promote" piece_reference piece_reference super_mask mover_arg? ")"
effect_set_score: "(" "set_score" mover_reference function ")"
@@
function_add: "(" "add" function+ ")"
function_connected: "(" "connected" piece_reference multi_mask_arg mover_arg? direction_arg? ")"
@@
function_count: "(" "count" super_mask ")"
function_line: "(" "line" piece_reference positive_int orientation_arg? exact_arg? player_arg? exclude_arg? ")"
function_multiply: "(" "multiply" function+ ")"
function_pattern: "(" "pattern" piece_reference (pattern_arg | shape_arg) rotate_arg? player_arg? exclude_arg? ")"
function_score: "(" "score" mover_reference ")"
function_subtract: "(" "subtract" function function ")"
@@
end_rules: "(" "end" end_rule+ ")"
end_rule: "(" "if" super_predicate end_rule_result ")"
@@
result_win: "(" (mover_reference | BOTH) "win" ")"
result_lose: "(" (mover_reference | BOTH) "lose" ")"
@@
super_mask: mask | mask_like_function | super_mask_and | super_mask_or | super_mask_not
super_mask_and: "(" "and" super_mask+ ")"
super_mask_or: "(" "or" super_mask+ ")"
super_mask_not: "(" "not" super_mask ")"
@@
| mask_promoted
| mask_row
| mask_region
mask_adjacent: "(" "adjacent" super_mask direction_arg? ")"
mask_captured: "(" "captured" ")"
mask_center: "(" "center" ")"
mask_column: "(" "column" nonnegative_int ")"
mask_corners: "(" "corners" ")"
mask_corner_custodial: "(" "corner_custodial" piece_reference mover_arg? ")"
mask_custodial: "(" "custodial" piece_reference custodial_length_arg mover_arg? orientation_arg? ")"
mask_edge: "(" "edge" (edge | FORWARD | BACKWARD) ")"
mask_empty: "(" "empty" ")"
mask_hopped: "(" "hopped" ")"
mask_occupied: "(" "occupied" mover_reference? ")"
mask_prev_move: "(" "prev_move" mover_reference ")"
mask_promoted: "(" "promoted" ")"
mask_row: "(" "row" nonnegative_int ")"
mask_region: "(" "region" region_reference ")"
// "Mask-like" functions are special functions that also support returning a mask. Currently this is only
// used for "line", but it might include other functions in the future
mask_like_function: function_line
@@
multi_mask_corners: "(" "corners" ")"
multi_mask_edges: "(" "edges" ")"
multi_mask_edges_no_corners: "(" "edgesNoCorners" ")"
@@
super_predicate_and: "(" "and" super_predicate+ ")"
super_predicate_or: "(" "or" super_predicate+ ")"
super_predicate_not: "(" "not" super_predicate ")"
predicate: predicate_action_was
| predicate_can_move_again
| predicate_equals
@@
predicate_action_was: "(" "action_was" mover_reference move_type_name ")"
predicate_can_move_again: "(" "can_move_again" move_type_name ")"
predicate_equals: "(" "=" function+ ")"
predicate_exists: "(" "exists" super_mask ")" // technically equivalent to (>= (count mask) 1)
@@
predicate_greater_equals: "(" ">=" function function ")"
predicate_last_move_in: "(" "last_move_in" super_mask ")" // equivalent to (exists (and (prev_move mover) mask)), but more readable
predicate_less_equals: "(" "<=" function function ")"
predicate_mover_is: "(" "mover_is" player_reference ")"
predicate_no_legal_actions: "(" "no_legal_actions" ")"
predicate_passed: "(" "passed" (mover_reference | BOTH) ")"
@@
direction_arg: "direction:" (direction | multi_direction)
distance_arg: "distance:" positive_int
@@
exclude_arg: "exclude:" multi_mask_arg
hop_over_arg: "hop_over:" (player_reference | mover_reference)
@@
mover_arg: "mover:" (mover_reference | BOTH)
multi_mask_arg: multi_mask | super_mask | "(" super_mask+ ")"
@@
pattern_arg: "(" positive_int indices_arg ")" // The positive int is the "width" of the pattern and the indices are the positions within the pattern
indices_arg: "(" nonnegative_int+ ")"
player_arg: "player:" (player_reference | mover_reference)
piece_arg: "piece:" piece_reference
priority_arg: "priority:" nonnegative_int
@@
same_piece_arg: "same_piece:" boolean
shape_arg: shape_definition
shape_definition: hexagon_shape
| hex_rectangle_shape
| rectangle_shape
| square_shape
hexagon_shape: "(" "hexagon" odd_int ")"
hex_rectangle_shape: "(" "hex_rectangle" positive_int positive_int ")"
rectangle_shape: "(" "rectangle" positive_int positive_int ")"
square_shape: "(" "square" positive_int ")"
// Optional rendering details
rendering: "(" "rendering" rendering_detail+ ")"
rendering_detail: color_assignment
| piece_shape_assignment
color_assignment: "(" "color" player_reference color ")"
piece_shape_assignment: "(" "shape" piece_reference piece_shape ")"
@@
?direction: true_direction | relative_direction
?move_type_name: HOP | SLIDE | STEP
?multi_direction: "(" direction+ ")"
?true_direction: UP | DOWN | LEFT | RIGHT | UP_LEFT | UP_RIGHT | DOWN_LEFT | DOWN_RIGHT | VERTICAL | HORIZONTAL | ORTHOGONAL | DIAGONAL | BACK_DIAGONAL | FORWARD_DIAGONAL | ANY
?relative_direction: FORWARD | BACKWARD | FORWARD_LEFT | FORWARD_RIGHT | BACKWARD_LEFT | BACKWARD_RIGHT
@@
?color: WHITE | BLACK
?piece_shape: CIRCLE | SQUARE | TRIANGLE | STAR | DIAMOND
@@
Appendix C Benchmark Game Descriptions
Below, we present natural language descriptions of the rules for each of the exemplar games analyzed in Section˜6.